OpenStudy (anonymous):

If sin(theta) = (1/2)x, what would sin(4 theta) equal?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Can you even solve it?

OpenStudy (anonymous):

cos(theta)=sqrt(1-x^2/4) so sin(2 theta)=2(x/2)sqrt(1-x^2/4) =xsqrt(1-x^2/4) cos(2 theta)=sqrt(1-x^2(1-x^2/4)) now find sin(4 theta) =2sin(2theta)cos(2theta)

OpenStudy (anonymous):

\[\sin(4 \theta ) = 2\sin(2 \theta ) \cos(2 \theta) = 2\sin(\theta)\cos(\theta) [ 1-2\sin^2 (\theta)] \]

OpenStudy (anonymous):

It should equal to 4 sin(theta) cos (theta) [ 1- 2 (sin ^ 2 (theta) )]